The following information relates to the debt securities investments of Sarasota Company.

1. On February 1, the company purchased 12% bonds of Gibbons Co. having a par value of $330,000 at 100 plus accrued interest. Interest is payable April 1 and October 1.

2. On April 1, semiannual interest is received.

3. On July 1, 10% bonds of Sampson, Inc. were purchased. These bonds with a par value of $181,200 were purchased at 100 plus accrued interest. Interest dates are June 1 and December 1.

4. On September 1, bonds with a par value of $57,600, purchased on February 1, are sold at 99 plus accrued interest.

5. On October 1, semiannual interest is received.

6. On December 1, semiannual interest is received.

7. On December 31, the fair value of the bonds purchased February 1 and July 1 are 95 and 93, respectively.

(a) Prepare any journal entries you consider necessary, including year-end entries (December 31), assuming these are available-for-sale securities. (Note to instructor: Some students may debit Interest Receivable at date of purchase instead of Interest Revenue. This procedure is correct, assuming that when the cash is received for the interest, an appropriate credit to Interest Receivable is recorded.) (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Date

Account Titles and Explanation

Debit

Credit

(1)

Feb. 1

(2)

(3)

Jul. 1

(4)

Sep. 1

(5)

(6)

(7)

(To record interest.)
Answer #1
No. Date Account Titles and Explanation Debit Credit
1 Feb 1 Debt Investment (available-for-sale securities) $      330,000
Interest Revenue (Accrued interest = Oct prior period to Jan = 4 Months) (330000*12%*4/12) $         13,200
Cash $      343,200
To record the purchased 12% bonds of Gibbons Co.
2 Apr 1 Cash (330000*12%*6/12) $         19,800
Interest Revenue $        19,800
To record the semiannual interest received.
3 July 1 Debt Investment (available-for-sale securities) $      181,200
Interest Revenue (Accrued interest = June = 1 Months) (181200*10%*1/12) $           1,510
Cash $      182,710
To record the purchased 10% bonds of Sampson, Inc.
4 Sep 1 Cash ((57600*99%)+2880) $         59,904
Loss on sale of debt investment (57600*1%) $               576
Debt Investment (available-for-sale securities) $        57,600
Interest Revenue (Apr to Aug = 5) (57600*12%*5/12) $          2,880
To record the sold 12% bonds of Gibbons Co.
5 Oct 1 Cash $         16,344
Interest Revenue (272400*12%*6/12) $        16,344
To record the semiannual interest received. (Remaining bonds of Gibbons = 330000-57600 = 272400
6 Dec 1 Cash $           9,060
Interest Revenue (181200*10%*6/12) $          9,060
To record the semiannual interest received.
7 Dec 31 Interest Receivable $         12,406
Interest Revenue (Notes 1) $        12,406
To record the accrue interest.
Dec 31 Unrealized Holding loss - OCI $         26,304
Fair Value Adjustment $        26,304
To record fair value adjustment.
Interest accrued on Gibbond (Sep to Dec = 4 months) (272400*12%*4/12) $        10,896
Interest accrued on Sampon (Dec = 1 month) (181200*10%*1/12) $          1,510
Accrue interest $        12,406
Available-for-Sale Portfolio
Security Cost Fair Value Unrealized Gain (Loss)
Gibbons Co. (Fair value = 272400*95%) $      272,400 $      258,780 $ (13,620)
Sampson, Inc. (Fair value = 181200*93%) $      181,200 $      168,516 $ (12,684)
Total $      453,600 $      427,296 $ (26,304)